The air above water is cooled as a result of the water absorbing heat from the air. How large a volume of air would be cooled by 1 degree Celsius if 1.21kg of water increases in temperature by 2.7 degree Celsius. The density of air is 1.29 kg/m 3 and the specific heat of air is 1000 J/kg oC.
Answer #1

Absorbed heat = 1.21 x 2.7 x 4.2 = 13.72 J

Mass of air, m = 13.72/1000 = 0.0137 kg

Volume, V = 0.0137/1.29 = 0.01064 m^3 = 10.64 litre

Comment in case any doubt please rate my answer....